OpenStudy (anonymous):

please help with this parametric equation: x= 3Sin^3t y= 3Cos^3t

OpenStudy (anonymous):

what you need help?

OpenStudy (anonymous):

how would i go about graphing these types of problems?

OpenStudy (anonymous):

 choose the regular values of "t" between \(-2\pi\le t\le2\pi\)